OpenStudy (anonymous):

Find the slope between points (2, 3) and (-1, 0).

hero (hero):

Insert points into y = mx + b 3 = 2m + b 0 = -m + b m = b 3 = 2m + m 3 = 3m 3/3 = m 1 = m The slope is 1

OpenStudy (anonymous):

Thank you!

hero (hero):

I hope you understand what was done to find the slope

OpenStudy (anonymous):

let first point be A(2 , 3) and other be B(-1 , 0) the slope of AB =(0 - 3)/(-1 -2) = -3/-3 =1